Circular No. 6294

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NGC 2363 L. Drissen, J.-R. Roy and C. Robert, Département de Physique, Uni-

versité Laval, Québec, communicate: “We report the discovery of an unusually bright new object in the middle of the giant H II region NGC 2363, based on Hubble Space Telescope WFPC2 images obtained on Jan.

8. Preliminary reductions indicate V = 17.95, B − V ∼ 0 for the new object. This point source, which was then the brightest star in NGC 2363 (M_{V} ∼ −10.0), was not visible on groundbased CCD images obtained in Jan. 1991 and Oct. 1992 at the Canada-France-Hawaii telescope; although crowding is severe in groundbased images, no point source brighter than V ∼ 22 was visible at that location. WFPC2 images indicate that this object is a strong Hα point source, surrounded by a small (radius 0′′.09, or 1.5 pc at a distance of 3.5 Mpc) Hα shell. We suspect that this object is a Luminous Blue Variable star in eruption (à la η Car). The coordinates of this star, from our WFPC2 images, are α = 7^{h}28^{m}43^{s}.4, δ = +69^{o}11′24′′ (equinox 2000.0). This is 0′′.5 west and 1′′.3 north of the ‘eastern knot’ of NGC 2363 (Fig. 2 in Drissen et al. 1993, A.J. 106, 1460). NGC 2363 is the brightest star-forming region in the Magellanic irregular NGC 2366, a member of the M81 group.”

3C 279 A. Wehrle, Infrared Processing and Analysis Center, California Insti-

tute of Technology; and R. C. Hartman, Goddard Space Flight Center. on behalf of the EGRET Team and Blazar Multiwavelength Consortium, report that 3C 279 is undergoing a γ-ray flare, as observed by EGRET on the Compton Gamma Ray Observatory (CGRO): “The EGRET quicklook analysis for Jan. 16–21 shows that 3C 279 was brighter in γ rays (energy E above 100 MeV) than in any observation since 1991, with a roughly estimated flux of 1.3 to 1.4 × 10^{−6} photon cm^{−2} s^{−1} at E > 100 MeV (probable error ∼ 20 percent). This is about a 10σ detection. The CGRO observation of 3C 279 is being extended by one week, to Feb. 6, to follow the development of this outburst. As part of an extensive multiwavelength campaign, observations with other spacecraft (IUE, XTE, ROSAT), as well as numerous groundbased observations, are currently underway or have been concluded recently. Additional groundbased observations are sought, especially at optical and near-infrared wavelengths. T. Balonek, Colgate University, reports V = 15.3, R = 14.8 on Jan. 18; variations of one or more magnitudes, on timescales of days to weeks, have been characteristic of this object during the past few years.”

1996 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden